OpenStudy (anonymous):

Choose all answers that are correct. 3. (1 pt) Which literal equations are equivalent to b=A/h ? Choose all answers that are correct. A.h=bA B.h=A/b C.A=bh D.A=b/h

OpenStudy (anonymous):

fan and medal for help!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

A and B

OpenStudy (anonymous):

thanx

OpenStudy (anonymous):

good question, let me remember math i haven't used in years

OpenStudy (anonymous):

Which literal equation is equivalent to 1.4 = kx A. k=x/1.4 B.x=1.4 x k C.k=1.4/x D.x=k/1.4

OpenStudy (anonymous):

C

OpenStudy (anonymous):

Solve for v. m=bv/k A. v = mb/k B. v=mkb C. v = b/mk D. v = mk/b

OpenStudy (anonymous):

Am I doing your homework for you?

OpenStudy (anonymous):

no definitely not

OpenStudy (anonymous):



OpenStudy (anonymous):

divide?

OpenStudy (anonymous):

same rule. use multiplication on both side to deal with division and vice versa

OpenStudy (sleepyjess):

@mattherd , I see you are new here. Welcome to OpenStudy! Please do not give out direct answers.